OpenStudy (anonymous):

a right triangle has a hypotenus of 13 and one leg 12cm.what is the other leg.

OpenStudy (anonymous):

anyone pleaseeeeeeeeee

OpenStudy (amistre64):

do you recall the pythag thrm?

OpenStudy (anonymous):

i did that and i got my answer wrong

OpenStudy (amistre64):

then you didnt "do that" :) leg^2 + 12^2 = 13^2 leg^2 = 169 - 144 leg^2=25

OpenStudy (anonymous):

thnak you

OpenStudy (anonymous):

thankx

OpenStudy (amistre64):

3,4,5 and 5,12,13 are common rt triangles that they like to give out in questions :)